OpenStudy (volkoner):

Consider the following trigonometric function: g(x)=−sin(8x−3)+5. What is the midline equation of the function? Give an exact expression.

Directrix (directrix):

Definition of Midline The midline is a horizontal axis that is used as the reference line about which the graph of a periodic function oscillates. Read more here; http://www.icoachmath.com/math_dictionary/midline.html @ http://www.icoachmath.com/math_dictionary/midline.html

Directrix (directrix):

@Volkoner If you are live in this thread now, we can do this problem now.

Directrix (directrix):

The equation of the midline of periodic function is the average of the maximum and minimum values of the function.

OpenStudy (volkoner):

Ok I will try to figure it out from the information you gave me

OpenStudy (volkoner):

So i understand the formula needed but i'm not sure how to get the numbers for parts of the formula

OpenStudy (volkoner):

@directrix the max is 5? I'm not sure what the amplitude is. 8?

Directrix (directrix):

The amplitude is the absolute value of (-1) −sin(8x−3)+5.

Directrix (directrix):

Follow this: The equation of the midline of periodic function is the average of the maximum and minimum values of the function.

Directrix (directrix):

What is the maximum value of sine of x? What is the minimum value of sine of x? Post those two and with a little more work, you'll have the midline equation.

Directrix (directrix):

Take a look at your graph.

OpenStudy (volkoner):

Thank you, I did 6+4 then divided that by 2 to get my answer. y=5

OpenStudy (volkoner):

I don't understand how you got the max and min though.

Directrix (directrix):

You could look on the graph but that is not necessary if you know that the maximum that sine of an angle can be in its life is 1 and the least that sine of an angle can ever be in its life is -1. Look at −sin(8x−3)+5 as Negative [ sine (angle)] + 5. Max value of Negative [ sine (angle)] + 5 will be Negative* (-1) + 5 = 6 Min value of Negative [ sine (angle)] + 5 will be Negative * (1) + 5 = -1 + 5 = 4.